OpenStudy (anonymous):

sin2x+cos2x=0 find all solution in 0 is less than or equal to x < 2pi

OpenStudy (freckles):

If it makes it less confusing write it as this: let u=2x (so u/2 =x) I'm going to put this strictly in terms of u like so sin(u)+cos(u)=0 0 <= u/2<2pi (solve for u) sin(u)=-cos(u) 0< =u<4pi so you need to find when sin(u) is the opp of -cos(u) for when u is between 0 and 4pi

OpenStudy (freckles):

use unit circle

OpenStudy (freckles):

http://etc.usf.edu/clipart/43200/43215/unit-circle7_43215_lg.gif here look at the order pairs when are the ordered pairs opposites (ex. opp of 3 is -3; -3 and 3 are opps)

OpenStudy (anonymous):

i dont get it

OpenStudy (freckles):

I'm asking you to look at the ordered pairs and tell for which angles they are opposite?

OpenStudy (freckles):

the only time you will have they are opposite is when one is sqrt(2)/2 and the other is -sqrt(2)/2